Gamma rays (γ-rays) are high-energy photons. In a certain nuclear reaction, a γ-ray of energy 0.505 MeV (million electronvolts) is produced. Compute the frequency of such a photon.
answer in Hz?
Answer #1

Solution) Energy , E = 0.505 MeV

1 MeV = 10^(6)×1.6×10^(-19) J

E = 0.505×10^(6)×1.6×10^(-19) J

E = 0.808×10^(-13) J

Frequency , f = ?

E = hf

Here h is plancks constant

h = 6.625×10^(-34) Js

E = hf

f = (E)/(h)

f = (0.808×10^(-13))/(6.625×10^(-34))

f = 1.219×10^(20) Hz

Approximately , f = 1.22×10^(20) Hz